For a two-tailed t-test with α = .30 and n = 125 the rejection zone is _____.
Answer #1

df = 125 - 1 = 124

0.3 / 2 = 0.15

critical value = + t0.15,124 = + 1.041

Rejection zone is t < -1.041 or t > 1.041